What is the temperature of 0.620 mol of gas at a pressure of 1.4 atm and a volume of 11.8 L ?
Answer #1

PV = nRT

where, P = pressure = 1.4 atm

V = volume = 11.8 L

n = number of moles = 0.620 mol

R = Gas constant

T = temperature = ?

1.4 * 11.8 = 0.620 * 0.0821 * T

16.52 = 0.0509 * T

T = 16.52 / 0.0509 = 325 K

T = 325 - 273 = 52 oC

Therefore, temperature = 52 oC